OpenStudy (anonymous):

Need help: Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 f(n) = 35n Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means. Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

Someone?

OpenStudy (anonymous):

1) written recursively f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 written explicitly f(n) = 35n 2) f(2)= f(1)+f(1) =35 + 35=2*35 f(3)= f(1)+f(2) =35 + 2*35=3*35 f(4)= f(1)+f(3) =35 + 3*35=4*35 f(5)= f(1)+f(4) =35 + 4*35=5*35 3) f(40)=40*35 = 1400 4) f(1)=5 ; f(n)= f(n-1) + (n-1) for n>1

OpenStudy (anonymous):

Thank you!!!